If a ray travelling along the line x = 1 gets reflected from the line x + y = 1, then the equation of the line along which the reflected ray travel is

Grade:12

1 Answers

Satyadip Mahapatra
8 years ago
the angle between x = 1 and x + y = 1 is 45 degrees.
as angle i = angle r
so, angle r = 45 degrees
putting the value in the formula  tan theta = (m1 – m/ 1 + m1m2 )
we get m = 0
so the answer wiil be y = 0 / y = c.